The plate with two narrow cracks at a distance of 0.4 mm was illuminated with monochromatic light of wavelength 633 nm. A light strip 9.5 mm from the center of the curtain was observed on a curtain 200 cm behind the panel. The order of the observed maximum is: a) 2 b) 3 c) 4 d) 5 e) 6


1
Expert's answer
2021-03-14T19:13:06-0400

"d\sin\Theta=m\lambda"


"\sin\Theta\approx\tan\Theta\approx\frac{x}{L}=\frac{0.0095}{2}=0.00475"


"0.0004\cdot0.0475=m\cdot633\cdot10^{-9}\to m=3"


"b)\ 3" . Answer